PLEASE HELP ME WITH THIS PROBLEM !!!=[
timmy is 4 years old.Janet is 3 times as old as timmy.when timmy turns 12,how old will janet be ?

Answer by stanbon(75887)

timmy is 4 years old.
t = 4
------------------
Janet is 3 times as old as timmy.
J = 3t = 3*4 = 12
-------
when timmy turns 12,how old will janet be ?
==================
Since J = 3t,
If timmy is 12, Janet will be 3*12 = 36
=========================================
Cheers,
Stan H.